OpenStudy (anonymous):

Solve by Factoring: x^4 - x^3 + x - 1=0

OpenStudy (anonymous):

Simplify the expression 3(21m + 27).

OpenStudy (anonymous):

you distribute the 3 to the 21m and the +27

OpenStudy (anonymous):

so it would be (3*21m) + (3*27)

OpenStudy (ghazi):

x^3(x-1)+(x-1)=0 (x^3+1)(x-1)=0

OpenStudy (anonymous):

how did you get that answer?

OpenStudy (ghazi):

just take x^3 common from x^4 +x^3

OpenStudy (anonymous):

ok and then you just keep the x+1?

OpenStudy (anonymous):

**x-1

OpenStudy (lopus):

 this is ruffinii